7. Mitchell decided to lease a car for his new job. He makes a down payment of $1700 and payments at the end of each month of $233 for 3 years.




 a) Write an equation to model the relationship between the total amount spent on the lease, A, and time, t.




 b) Use the equation to determine what Mitchel will have paid by the end of the 1" year, and by the end of the 3rd year.





 C) Draw a graph of the equation would look like for 0 to 4 years. What do you notice?






 d) Write an equation to model only the amount spent in monthly payments over three years. How do you know this relationship is no longer a partial variation?





 e) Another dealership offers Mitchell the same car, but with $1000 down and monthly payments of $272 for 3 years. Which lease should Mitchel choose? Why?



1
Expert's answer
2021-05-07T09:09:17-0400

a.) A=233t + 1700

b.) "A=233\times12+1700=3,496"

"A=233\times36+1700=7,059"

c.) Drawn




d.) Equation no longer a partial equation because time is used up and all amount repaid.

e.) "A=272\times36+1000=10,792"

Use first one to save 3,733